Point B is between A and C. AB is three times BC. If AB = 81, find AC.
If possible please show your work,
Thanks.

Answer by rothauserc(4718)

AB = 3 * BC
81 = 3 * BC
divide both sides of = by 3
BC = 27
AC = 27 + 81
AC = 108